OpenStudy (konradzuse):

solve the differential eq y' = 53x * e^x^2 y = ?

OpenStudy (konradzuse):

\[y' = 53x * e^{x^2}\]

OpenStudy (anonymous):

\[\frac{dy}{dx}=53xe^ {2x}\]

OpenStudy (konradzuse):

2 comes down?

OpenStudy (konradzuse):

these diff eq's always konfuse me...

OpenStudy (konradzuse):

Shouldn't I just solve normally by parts?

OpenStudy (anonymous):

are you in differential equations

OpenStudy (anonymous):

or is this a calculus question

OpenStudy (konradzuse):

calc 2.

OpenStudy (anonymous):

well first multiply by dx

OpenStudy (anonymous):

\[dy=e^{x^2}53xdx\]

OpenStudy (anonymous):

if you integrate both sides

OpenStudy (anonymous):

you get y=

OpenStudy (konradzuse):

oh yeah that's right... :P

OpenStudy (anonymous):

let u x^2 du=2xdx du/2=xdx

OpenStudy (konradzuse):

so hmm. u = 53x du = 53dx \[dv = e^{x^2}\] v = hmm

OpenStudy (anonymous):

\[y=53/2\int e^u\]

OpenStudy (konradzuse):

ah...

OpenStudy (anonymous):

no need by parts only ned u sub

OpenStudy (konradzuse):

okies, I'm thinking too much with mega headache :P.

OpenStudy (konradzuse):

\[y = \frac{53}{2} \frac{e^{2u}}{2}\]

OpenStudy (konradzuse):

or would it be... \[e^{u^2}\]

OpenStudy (konradzuse):

oosp nvm it stays the same...

OpenStudy (konradzuse):

\[y = \frac{53}{2} * e^u + c\]

OpenStudy (konradzuse):

\[y = \frac{53}{2} * e^{x^2} + c\]

OpenStudy (anonymous):

yep